OpenStudy (nyquasha13):

Express each product in the simplest form. 4 . 4k - - k 4 A. 1 B. 4 C. 16k --- 4 D. 4k

OpenStudy (astrophysics):

\[\frac{ 4 }{ k } \times \frac{ 3k }{ 4 } ?\]

OpenStudy (astrophysics):



OpenStudy (nyquasha13):

change that 3k to a 4k

OpenStudy (astrophysics):